OpenStudy (anonymous):

Determine the present value P you must invest to have the future value A at simple interest rate r after time t. Assume there are 52 weeks in a year and round to the next cent. A = $5,000.00, r = 9.0%, t = 39 weeks A=P(1+rt)

OpenStudy (anonymous):

P = P/ ( 1 + r * t ) Can you plug the number into your calculator?

OpenStudy (anonymous):

yeah I can. Basically i didn't know how to convert 39 weeks into T

OpenStudy (anonymous):

t = 39/ 52 = 3/4

OpenStudy (anonymous):

thanks! For some reason I was getting .82

OpenStudy (anonymous):

Divide by 13!

OpenStudy (anonymous):

So you're fine from here ?

OpenStudy (anonymous):

I think so. Here's my formula 5,000= P(1+9*.75)

OpenStudy (anonymous):

lol you've come to my rescue again!

OpenStudy (anonymous):

sh*t I did that wrong.

OpenStudy (anonymous):

it was supposed to be .09 not 9

OpenStudy (anonymous):

yep, I'm just paying attention to your formula , forget the number:

OpenStudy (anonymous):

P = $ 4,683.84

OpenStudy (anonymous):

Are you okie now?

OpenStudy (anonymous):

yeah...well I didn't pass my test, so I have to redue it. :( will college ever end!

OpenStudy (anonymous):

When did you have the test? what test is it?

OpenStudy (anonymous):

I'm taking it online. It's an online math course

OpenStudy (anonymous):

I'm sorry that you didn't pass, however, frankly you should invest more effort into your study!

OpenStudy (anonymous):

lol as much as I agree with you. and I do. Working 40 hours and having to clean a house, take care of a dog,finding a big girl job, working out...well It takes up a lot of time. I basically have from 9 to 1am to study, and do h/w.

OpenStudy (anonymous):

If so, that's totally understandable! How could you survive with full time job and college work load? What do you mean "finding a big girl job" ?

OpenStudy (anonymous):

it's a term I use for a grown up job, like a not minimum wage- career starting job. haha, idk. So needless to say applications take time, but I don't really have a choice but to work full time in a not big girl job to pay for bills, school, etc!

OpenStudy (anonymous):

You're so brave! I just volunteer in weekend and I'm already out of breath!

OpenStudy (anonymous):

Anyway, just message me whenever you need help :)

OpenStudy (anonymous):

haha thanks for your help! I will be in desperate need of any help between now and Wednesday.

OpenStudy (anonymous):

All you do is just post them up :)

OpenStudy (anonymous):

lol i try to figure the ones i can out :) i'm not a complete slacker! haha